28. Find the area of the indicated region under the standard normal curve. A normal curve is over a horizontal axis and is centered on 0. Vertical line segments extend from the horizontal axis to the curve at negative 0.45 and 2.11. The area under the curve between negative 0.45 and 2.11 is shaded. A. 1.3090 B. 0.3438 C. 0.6562 D. 0.3090
Answer #1

28)

Solution :

P(-0.45 < z < 2.11) = P(z < 2.11) - P(z < -0.45) = 0.9826 - 0.3264 = 0.6562

C) Area = 0.6562